OpenStudy (anonymous):

In this lesson, you learned how to transform square-root functions. Transformations may be: reflection (across the x- or y-axis); translation (horizontal or vertical); stretch (horizontal or vertical); or compression (horizontal or vertical). Suppose you start with a parent function f(x) = √x. Which of the four transformations is easiest for you to represent and which is the most difficult? Why? Give an example of a "child" of the parent function. Make sure you identify the transformation so your reader knows exactly what happened and include the words vertical or horizontal, if necessary.

OpenStudy (here_to_help15):

Let us take a function y=f(x). assuming limit of function exist (rectangular coordinate) 1. Horizontal shifting: y=f(x-a) where a is constant this shifts curve to the the left by a units y=f(x+a) this shift curve to the right by a units. 2. Vertical shifting: y=f(x)+a shift curve to the negative y axis by a unit y=f(x)-a shift curve to the positive y axis by a unit. 3.horizontal stretching take a non zero quantity y=f(x/a) this stretch the curve in horizontal direction 4. horizontal compressing y= f(ax) this compress curve in horizontal by a factor 5. vertical stretching y=af(x) this stretch curve in vertical 6. vertical compressing' y=f(x)/a this compress curve in vertical 7.y=x line gives the inverse of the function for inverse just replace x by y

OpenStudy (here_to_help15):

I hope this helps : )